# Learning in Structured Stackelberg Games

### v1

@@ 3.1.3 Relationship with Littlestone dimension
We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting. We next explore how that relates to multiclass learnability. While, as discussed, learning to predict the follower type from the context is a sufficient condition for learning the leader’s optimal policy, in this section we prove that this is not a necessary condition. In Theorem 3.11, we show that the Littlestone dimension is always at most as large as the Stackelberg Littlestone dimension and in Proposition LABEL:prop:ldim=sldim we show that equality always holds in the special case of two follower types. In Theorem 3.12, we show that an out-of-the-box application of the optimal algorithm for multiclass learning in our setting can result not only in suboptimal number of mistakes, but also in suboptimal utility for the leader. This is in contrast to our proposed algorithm, which achieves optimal utility by accounting for the interaction between the hypothesis class and the Stackelberg game.
Theorem 3.11.
@@
For any hypothesis class $\pazocal{H}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
Proof.
If the leader plays SOA, she will make at most $\textnormal{Ldim}(\pazocal{H})$ mistakes. This is because she will predict the follower’s type incorrectly at most $\textnormal{Ldim}(\pazocal{H})$ times, and predicting the follower’s type correctly is a sufficient condition for playing the optimal mixed strategy. Since the SOA makes $\textnormal{Ldim}(\pazocal{H})$ mistakes and we know that any deterministic algorithm will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by Theorem 3.8, we can infer that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$. ∎
@@
While it is known that SOA makes at least $\textnormal{Ldim}(\pazocal{H})$ mistakes when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that SOA achieves suboptimal regret in our setting. This is because mispredicting the follower type does not necessarily imply that the optimal strategy against that and the true follower type do not yield the same utility for the leader. In Theorem 3.12, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
@@
Consider $\pazocal{Z}=\{1,2,3,4\}$ and $\mathscr{K}=\{f^{(1)},f^{(2)},f^{(3)}\}$. We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
@@
Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure 2.
Figure 2: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a labeling of the parent node to a follower type.
Consider now a Stackelberg game with utility functions $u^{(1)},u^{(2)},u^{(3)}$ corresponding to follower types $1,2$, and $3$ respectively, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$. Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$. Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table 1.
@@
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$. We will consider any possible SL tree. First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered SL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$). So it suffices to consider contexts $1$ and $2$. In Figure 3 we show the two $\pazocal{H}$-shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively. Notice that the maximal depth of a full binary tree is $1$, as desired.
Say that the adversary presents context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each. To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$. Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree. The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure 2. So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$. However, the true follower type is $f_{1}^{*}=f^{(1)}$ (mistake).
  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$. No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks. The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm. So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (mistake).
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is $1$ and we have shown that for the above sequence of contexts SOA makes $2$ mistakes, this concludes our proof.
For ease of exposition, we also provide an overview of SSOA for the same series of nodes (see Figure (3(a)) for reference).
  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$. Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth $1$, while the second branch cannot lead to a full binary tree. So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no mistake).
  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree. SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by SSOA (mistake).
@@
While we have shown that running multiclass SOA alone can in certain settings achieve suboptimal utility for the leader, in the following theorem we show that our algorithm may be interpreted as running SOA in a transformed space.
Theorem 3.13.
For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes. Moreover, this algorithm coincides with SSOA.
Proof.
Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that:
$\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
Notice that $\zeta$ is a bijection given that each follower type is distinct (see Remark 2.1). Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$. This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\max_{\mathbf{x}\in\Delta(A)}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (see Theorem 1 in [22]). Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Consider any SL-Tree for $\pazocal{H}$. Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well. Also, for any root-to-leaf path on this tree that is consistent with a function $h\in\pazocal{H}$, there exists $\pi_{h}\in\Pi$ that is consistent with it as well, by definition of $\pi_{h}$. The converse is also true: in any Littlestone tree for $\Pi$ consider any root-to-leaf path, that is consistent with some $\pi_{h}\in\Pi$. Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it. Therefore, a tree is SL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$. This holds for all trees, so it will hold for the maximal-depth trees as well.
3.2 No-regret learning in the agnostic setting
Definition 3.14 ((Agnostic Setting) Contextual Stackelberg Regret).
where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
Combining our SSOA algorithm with results and tools from multiclass prediction [17], we get the following regret bound.
Theorem 3.15.
Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$. In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
Theorem 3.16.
Let $\pazocal{H}$ be any hypothesis class. There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite Stackelberg-Littlestone dimension $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
We include the complete proofs of Theorems 3.15 and 3.16 in Appendix A.3.
4 Distributional Learning
We now shift our focus from online learning to distributional learning. Suppose that there exists an unknown, joint distribution over contexts and follower types. We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function. When contexts come into play in Section 4.2, we analyze an intuitive approach that uses a black box empirical risk minimizer (ERM) and show that it is suboptimal in our setting. Analogous to our online results, we then give a better sample complexity bound that depends on a new complexity measure.
@@ 4.1 Learning without Contexts
Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S$ of follower types, we write:
$u_{S}(\mathbf{x})=\frac{1}{|S|}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$
Given a set $S$ of $m$ samples from $\mathscr{D}$, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$. In other words, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that with high probability:
$\mathbb{P}_{S}[\max_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon]\leq\delta.$
For this, it suffices to show a generalization guarantee, i.e. a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
@@
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta$
The implication of a generalization guarantee is that it relates the utility of the leader’s strategy that maximizes the average payoff over the training data to the expected utility of the optimal strategy. Formally, this means that for $\hat{x}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $x^{*}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta.$
Consider the class $\pazocal{P}=\{u_{\mathbf{x}}:\mathscr{K}\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader, parameterized by a follower type that take as input the leader’s mixed strategy and output the leader’s utility. Consider the following definition, introduced by [6] and adapted to our setting:
Definition 4.2.
A class $\pazocal{H}$ of hypotheses, $h:\mathscr{K}\to\mathbb{R}$, is $(d,t)$-delineable if:
  1. 1.
    The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$; and
  2. 2.
    For any $f\in\mathscr{K}$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
With this definition, we make the following useful proposition.
Proposition 4.3.
@@
The statement holds from the following two facts: (1.) Given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piece-wise linear function. (2.) For any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it. Each hyperplane corresponds to a change in best-response strategies for the follower from one pure action to another, so there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes. ∎

### v2

@@ 3.5 Relationship with Littlestone dimension
We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting. We next explore how that relates to multiclass learnability. While, as discussed, learning to predict the follower type from the context is sufficient to learn the leader’s optimal policy, in this section we prove that this is not a necessary condition. We start by showing that the Littlestone dimension upper bounds the Stackelberg Littlestone dimension. The proof of Lemma˜3.10 can be found in Section˜A.4.
Lemma 3.10.
@@
For any hypothesis class $\pazocal{H}$ and Stackelberg game, $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
@@
While it is known that any learner makes at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that the optimal multiclass learning algorithm (SOA) [22] achieves suboptimal regret in our setting. This is because mispredicting the follower type does not necessarily imply that the optimal strategy against the mispredicted and the true follower type do not yield the same utility for the leader. An illustrative example of this is our game construction in Theorem˜3.5, which shows that any learner can mispredict the follower type ad infinitum, yet the leader will achieve optimal cumulative utility regardless. In Theorem˜3.11, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm. [^5] In fact, there are cases in which SOA yields strictly lower utility even against the algorithm that minimizes policy mispredictions. We prove this result in Lemma B.6.
@@
[…]
@@
Notice that the Littlestone dimension of this hypothesis class is at least 222. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜1.
Figure 1: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a follower type.
Consider now a Stackelberg game with utility functions $u_{f^{(1)}},u_{f^{(2)}},u_{f^{(3)}}$, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$. Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$. Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table˜2.
@@
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$. We will consider any possible Stackelberg Littlestone tree. First, note that without loss of generality we can consider trees that do not include the context $z=4$ since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, for $z=3$, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
So it suffices to consider contexts 111 and 222. In Figure˜2 we show the two shattered Stackelberg-Littlestone tree constructions with $z=1$ and $z=2$ as the root nodes respectively. We start with the tree in Figure˜2(a). For $z=2$, its not hard to show that $\inf_{\mathbf{x}}\max_{j\in\{1,3\}}r(\mathbf{x},2,f^{(j)})=\inf_{\mathbf{x}}\max_{j\in\{2,3\}}r(\mathbf{x},2,f^{(j)})=1/2$, achieved at $\mathbf{x}=(1/2,1/2)$. For the root node $z=1$, we have that $\inf_{\mathbf{x}}\max\{r(\mathbf{x},1,f^{(1)})+1/2,\;r(\mathbf{x},1,f^{(2)})+1/2,\;r(\mathbf{x},1,f^{(3)})\}=3/4$ achieved at $\mathbf{x}=(3/4,1/4)$. Therefore, the weight of the root node is 3/43/43 / 4. A similar analysis shows that the root node weight of the SL-tree in Figure˜2(b) is 3/43/43 / 4. Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals 3/43/43 / 4.
We consider the case, where the adversary present context 111 followed by context 222. We go through the SOA outputs per timestep below.
[…]
  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$. No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks. The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm. So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (instantaneous regret equals 1).
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is 3/43/43 / 4 and we have shown that for the above sequence of contexts running SOA yields regret 222, this concludes our proof.
@@
4 Distributional Learning in Structured Stackelberg Games
We now shift our focus from online learning to distributional learning. Consider an unknown, joint distribution over contexts and follower types, from which the leader receives samples drawn i.i.d. The leader’s objective is to design her strategy in such a way to maximize the utility she will get on a new sample drawn from the same distribution. We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function. When contexts come into play in Section˜4.2, we prove necessary and sufficient conditions for PAC learnability.
@@ 4.1 Learning without Contexts
Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$. Concretely, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that
$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
To achieve this, it suffices to show a generalization guarantee i.e., a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
Definition 4.1 (Generalization guarantee).
$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
This implies a relationship between the utility of the leader’s strategy maximizing the average payoff over the training data to the expected utility of the optimal strategy. Formally, for $\hat{\mathbf{x}}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$
$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
We let $\pazocal{P}=\{u_{\mathbf{x}}:[K]\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$. Consider the following definition, introduced by Balcan et al. [9] and adapted to our setting.
@@
[…]
  • The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$,
  • For any $f\in[K]$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
We can now show that the primal class $\pazocal{P}$ is deliniable, which is a property we will use to show a generalization guarantee.
Lemma 4.3.
@@
The statement holds from the following two facts. First, given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piecewise linear function. Second, for any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it. Each hyperplane corresponds to a change in the follower’s best-response strategies, from one pure action to another. Therefore, there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes. ∎